Solve the following equations:
– 4 = 5(p – 2)..........

Grade:12

2 Answers

Harshit Singh
2 years ago
Dear Student

Let us divide both the side by 5,
= - 4/5 = (5(p–2))/5
= - 4/5 = p -2
By transposing - 2 from RHS to LHS it becomes 2
= - (4/5) + 2 = p
= (- 4 + 10)/ 5 = p
= p = 6/5

Thanks
Shraddha Singh
2 years ago
Dear Student,
The following question can be solved as follows :
 
-4 = 5(p-2)
-4 = 5p - 10
5p = -4+10
5p = 6
P = 6/5
 
Hope it helps!
Thank you ✌️